# Synergizing Motion and Appearance: Multi-Scale Compensatory Codebooks for Talking Head Video Generation

### v1

@@ 3.3 Joint Optimization Objective of the Framework
We use the multi-scale compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$ and the image decoder $\mathcal{D}_{I}$ for image decoding. The low-resolution feature $\mathbf{F}_{c}^{1}$ is fed as the initial input to $\mathcal{D}_{I}$, which gradually upsamples it through a series of upsampling layers and ResNet blocks (He et al., 2016) until it reaches the output resolution. The intermediate features are refined with higher-resolution features $\{\mathbf{F}_{c}^{i}\}_{i=2}^{N}$ by SFT (Wang et al., 2018) and addition when they match the resolution of $\mathbf{F}_{c}^{i}$. After fusing all the scales, $\mathcal{D}_{I}$ generates the final output $\mathbf{I}_{g}$.
The training objective combines the codebook losses from Sec. 3.2 with common losses for talking head video generation (Siarohin et al., 2019b; Hong et al., 2022). Specifically, we use the equivariance loss $\mathcal{L}_{eq}$ and keypoint distance loss $\mathcal{L}_{kpd}$ to guide keypoint prediction in the motion flow estimator, and an image reconstruction loss $\mathcal{L}_{recon}$ and an adversarial loss $\mathcal{L}_{adv}$ on $\mathbf{I}_{g}$. To avoid the image decoder relying too much on high-resolution appearance features, we also generate an image $\mathbf{I}_{g}^{1}$ using only $\mathbf{F}_{c}^{1}$ and apply $\mathcal{L}_{recon}$ to minimize the difference between $\mathbf{I}_{g}^{1}$ and $\mathbf{I}_{d}$. The overall training objective is:
$\mathcal{L}=\mathcal{L}_{recon}(\mathbf{I}_{d},\mathbf{I}_{g})+\lambda_{adv}\mathcal{L}_{adv}(\mathbf{I}_{d},\mathbf{I}_{g})+\mathcal{L}_{eq}+\mathcal{L}_{kpd}+\mathcal{L}_{vq,m}+\mathcal{L}_{vq,a}+\lambda^{1}\mathcal{L}_{recon}(\mathbf{I}_{d},\mathbf{I}_{g}^{1}),$ (6)
where $\lambda_{adv}$ and $\lambda^{1}$ are the corresponding loss term weights.
@@ 4.1 Implementation Details
Datasets. We conduct experiments on VoxCeleb1 (Nagrani et al., 2017) and CelebV-HQ (Zhu et al., 2022) datasets. We train our model on VoxCeleb1 training set. For evaluation, we build the test set on VoxCeleb1 by randomly sample 50 videos from its test split. To evaluate the model’s generalization ability, we randomly select 50 videos from CelebV-HQ for testing.
Metrics. For same-identity reconstruction, we adopt PSNR, $\mathcal{L}_{1}$ and LPIPS following (Tao et al., 2024) to evaluate the reconstruction quality. We also use FID (Heusel et al., 2017) to measure the realism of the generated video frames. Following (Siarohin et al., 2019a), we employ Average Keypoint Distance (AKD) for motion transfer quality evaluation and Average Euclidean Distance (AED) for identity preservation quality evaluation.
@@ 4.2 Comparison with State-of-the-art Methods
Figure 6: Qualitative comparison with state-of-the-art methods for (a) same-identity reconstruction and (b) cross-identity reenactment on the VoxCeleb1 and CelebV-HQ datasets. Our method can generate high-quality facial images under even large pose variations.
Table 1: Quantitative comparison with state-of-the-art methods for same-identity reconstruction on VoxCeleb1 and CelebV-HQ dataset. Our results are the best on the Voxceleb1 dataset and competitive on the CelebV-HQ dataset.
@@
FOMM (Siarohin et al., 2019b) 53.97 22.96 0.0474 0.2200 1.4037 0.1509 78.15 20.92 0.0685 0.2925 3.6098 0.2955
DaGAN (Hong et al., 2022) 51.55 22.92 0.0492 0.2251 1.5740 0.1652 99.84 20.49 0.0781 0.3209 7.6075 0.3301
MCNet (Hong & Xu, 2023) 51.45 24.59 0.0402 0.1996 1.2363 0.1254 78.33 22.20 0.0640 0.2732 4.1386 0.2903
MRFA (Tao et al., 2024) 48.49 25.26 0.0370 0.1872 1.1823 0.1188 75.73 22.41 0.0625 0.2670 3.7166 0.2527
AniPortrait (Wei et al., 2024) 52.65 20.15 0.0637 0.2767 2.6543 0.2623 61.56 19.71 0.0748 0.2878 2.2397 0.2739
FYE (Ma et al., 2024) 43.25 19.54 0.0714 0.2954 2.7071 0.2652 62.55 19.58 0.0802 0.3006 4.8637 0.3029
@@
We compare our method with a series of open-source state-of-the-art methods including non-diffusion based FOMM (Siarohin et al., 2019b), DaGAN (Hong et al., 2022), MCNet (Hong & Xu, 2023) and MRFA (Tao et al., 2024), and diffusion-based AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024).
Same-identity Reconstruction. To evaluate the performance on same-identity reconstruction, we use the first frame of each video as the source image and reconstruct the whole video. Tab. 1 presents the quantitative results for same-identity reconstruction. Our method outperforms the other methods almost on all metrics on VoxCeleb1 dataset and remains competitive when generalized to the more challenging CelebV-HQ. Compared with those unsupervised methods (Siarohin et al., 2019b; Hong et al., 2022; Hong & Xu, 2023; Tao et al., 2024), our method achieves better motion estimation, e.g., our method get the best AKD scores on CelebV-HQ dataset. This result verifies the effectiveness of our designed multi-scale motion codebook and its generalizability. For the results of image quality (i.e., FID, PSNR, $\mathcal{L}_{1}$, LPIPS), our method outperforms other methods on VoxCeleb1 dataset, even those diffusion methods (Wei et al., 2024; Ma et al., 2024) trained with larger-scale datasets. It indicates that our designed multi-scale appearance codebook is capable to compensate for the intermediate warped feature for better talking head video generation.We also show some qualitative results in Fig. 6a. Our method can generate plausible unseen facial regions (row 2) and handle large motion (row 4) with accuracy.
Cross-identity Reenactment. We conduct cross-identity reenactment experiments to validate our method. The qualitative results shown in Fig. 6b indicate the superiority of our method. Compared to non-diffusion-based MCNet (Hong & Xu, 2023) and MRFA (Tao et al., 2024), our method better preserves facial details (row 1 and 3), even under large motions (row 2 and 4). Diffusion-based FYE (Ma et al., 2024) often produced exaggerated expressions and struggled to imitate the driving expressions accurately, likely due to its reliance on landmark-based embeddings without explicitly modeling motion. These findings confirm the effectiveness of our framework. More experimental results are shown in Sec. A.2.2.
@@ 4.3 Ablation Study
We perform ablation studies to assess the effectiveness of the learned multi-scale motion and appearance compensatory codebooks. The model variants in Tab. 2 are as follows: (i) “Baseline” is the model without any compensatory codebook. (ii) “Baseline+SMC” includes only a single-scale motion codebook, compensating the initial motion flow only at scale $1$, and the compensated result is used to warp multi-scale features. (iii) “Baseline+MMC” includes a multi-scale motion codebook to compensate for the motion flows across scales. (iv) “Baseline+MMC+SAC” adds a single-scale appearance codebook to (iii), compensating only the warped feature at scale $1$. (v) “Baseline+MMC+MAC” is our full model with both multi-scale motion and appearance codebooks. We present the quantitative results in Tab. 2 and qualitative comparisons of (i), (iii), and (v) in Fig. 9
Effect of Joint Codebook Learning. To evaluate the effectiveness of our jointly learned multi-scale motion and appearance codebooks, we visualize the reconstructed multi-scale motion flows with the motion codebook in Fig. 7 and appearance features with the appearance codebook in Fig. 8. For the motion flows, results in Fig. 7 are output of the motion flow decoder $\mathcal{D}_{M}$ given the quantized motion flow features. Despite the limited number of codes, our multi-scale motion codebook can reconstruct high-quality motion flows, confirming its ability to capture typical local motion patterns. For appearance features, we visualize the quantized features directly in Fig. 8. Despite some quantization loss, the multi-scale appearance codebook reconstructs the driving features well with limited codes, demonstrating the its ability to store informative local appearance details. These results demonstrate the effectiveness of joint codebook learning, allowing the codebooks to store expressive local motion and appearance patterns.
@@
Figure 10: Visualization of the motion flow compensation process. We present the initial motion flow $\mathbf{M}^{0}$, the motion flow residual $\{\mathbf{M}_{r}^{i}\}_{i=1}^{N}$ and the compensated motion flows $\{\mathbf{M}^{i}\}_{i=1}^{N}$.
Figure 11: Visualization of the appearance compensation results. We present the warped source features $\{\mathbf{F}_{w}^{i}\}_{i=1}^{N}$ and the compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$.
Effect of Multi-scale Appearance Codebook Compensation. The fourth row in Tab. 2 shows that adding single-scale appearance codebook compensation on top of multi-scale motion codebook compensation improves FID, LPIPS, and AKD, indicating that appearance codebook compensation refines warped source features for more realistic image generation. However, there is a slight drop in PSNR, $\mathcal{L}_{1}$, and AED, likely due to a conflict between scale $1$ compensated appearance features and other warped features with warping artifacts. The fifth row shows consistent improvement, suggesting multi-scale appearance codebook compensation resolves this conflict and further refines warped features across scales, boosting overall performance. The last two columns in Fig. 9 also verify that multi-scale appearance codebook compensation leads to more accurate motion (e.g., the mouth in row 1, eyes in row 2, and shoulders in row 3) with realistic facial details (e.g., the hair in row 2). Additionally, Fig. 11 visualizes the compensated feature maps at different scales, showing more complete facial shapes and details compared to the warped feature $\mathbf{F}_{w}^{i}$. These results validate the effectiveness of multi-scale appearance codebook compensation.
@@ 5 Conclusion
Appendix A Appendix
A.1 Additional Implementation Details
@@
We perform multi-scale compensation across $N=4$ scales. We employ the keypoint-based motion flow estimator from FOMM (Siarohin et al., 2019b). The multi-scale motion flows are estimated at a size of $64\times 64$. We use convolution layers to encode the motion flows into a latent motion flow space of size $32\times 32\times 32$ and set the multi-scale motion codebook size to $K=1024$ and $d_{m}=32$. We also use convolution layers to decode the quantized motion flow features while adopting the motion flow updater in MRFA (Tao et al., 2024) as our motion flow residual decoder. We employ the image encoder and decoder architecture from VQGAN (Esser et al., 2021) and further encode the multi-scale appearance features into a size of $32\times 32\times 256$. The multi-scale appearance codebook size is set to $T=1024$ and $d_{a}=256$.
@@
We follow the unsupervised training pipeline from FOMM (Siarohin et al., 2019b), where the source and driving frames are extracted from the same video, and our framework learns to reconstruct the driving frame. For the training objective, we use the perceptual loss from FOMM (Siarohin et al., 2019b) along with the L1 loss as the image reconstruction loss, and we set the loss weights as $\lambda_{adv}=0.8$, $\lambda^{1}=0.5$, $\lambda_{recon,m}=32$ and $\beta=0.25$. We train the entire framework end-to-end using the Adam optimizer with a learning rate of $8\times 10^{-5}$ and a batch size of 16 for 250K iterations on four NVIDIA RTX 3090 GPUs.
A.2 More Details on Experiments
A.2.1 Additional Details on the Compared Methods
@@
We evaluate the performance of the compared methods using their released pre-trained models, and we present the training datasets used for each method in Tab. 3. All the GAN-based methods (Siarohin et al., 2019b; Hong et al., 2022; Hong & Xu, 2023; Tao et al., 2024) and our method are trained on the VoxCeleb1 (Nagrani et al., 2017) training set, while the diffusion-based methods AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024) are trained on larger-scale datasets, including VFHQ (Xie et al., 2022), CelebV-HQ (Zhu et al., 2022), HDTF (Zhang et al., 2021), and their own collected dataset (Ma et al., 2024).
A.2.2 More Experimental Results
Video Demo. We present a video demo along with the supplementary material, which includes video results from the ablation study and comparisons with state-of-the-art methods to demonstrate the performance of our video generation method.

### v2

@@ 3.3 Joint Optimization Objective of the Framework
We use the multi-scale compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$ and the image decoder $\mathcal{D}_{I}$ for image decoding. The low-resolution feature $\mathbf{F}_{c}^{1}$ is fed as the initial input to $\mathcal{D}_{I}$, which gradually upsamples it through a series of upsampling layers and ResNet blocks [13] until it reaches the output resolution. The intermediate features are refined with higher-resolution features $\{\mathbf{F}_{c}^{i}\}_{i=2}^{N}$ by SFT [30] and addition when they match the resolution of $\mathbf{F}_{c}^{i}$. After fusing features from all scales, $\mathcal{D}_{I}$ generates the final output $\mathbf{I}_{g}$.
The training objective combines the codebook losses from Sec. 3.2 with common losses for talking head video generation [24, 16]. Specifically, we use the equivariance loss $\mathcal{L}_{eq}$ and keypoint distance loss $\mathcal{L}_{kpd}$ to guide keypoint prediction in the motion flow estimator, and an image reconstruction loss $\mathcal{L}_{recon}$ and an adversarial loss $\mathcal{L}_{adv}$ on $\mathbf{I}_{g}$. To avoid the image decoder relying too much on high-resolution appearance features, we also generate an image $\mathbf{I}_{g}^{1}$ using only $\mathbf{F}_{c}^{1}$ and apply $\mathcal{L}_{recon}$ to minimize the difference between $\mathbf{I}_{g}^{1}$ and $\mathbf{I}_{d}$. The overall training objective is:
$\begin{split}\mathcal{L}=&\mathcal{L}_{recon}(\mathbf{I}_{d},\mathbf{I}_{g})+\lambda_{adv}\mathcal{L}_{adv}(\mathbf{I}_{d},\mathbf{I}_{g})+\mathcal{L}_{eq}+\mathcal{L}_{kpd}\\ &+\mathcal{L}_{vq,m}+\mathcal{L}_{vq,a}+\lambda^{1}\mathcal{L}_{recon}(\mathbf{I}_{d},\mathbf{I}_{g}^{1}),\end{split}$
where $\lambda_{adv}$ and $\lambda^{1}$ are the loss term weights.
@@ 4.1 Implementation Details
Datasets. We conduct experiments on VoxCeleb1 [20] and CelebV-HQ [46] datasets. We train our model on VoxCeleb1 training set. For evaluation, we build the test set on VoxCeleb1 by randomly sample 50 videos from its test split. To evaluate the model’s generalization ability, we randomly select 50 videos from CelebV-HQ for testing.
Metrics. For same-identity reconstruction, we adopt PSNR, $\mathcal{L}_{1}$ and LPIPS following [25] to evaluate the reconstruction quality. We also use FID [14] to measure the realism of the generated video frames. Following [23], we employ Average Keypoint Distance (AKD) for motion transfer quality evaluation and Average Euclidean Distance (AED) for identity preservation quality evaluation.
@@ 4.2 Comparison with State-of-the-art Methods
Figure 6: Qualitative comparison with state-of-the-art methods for (a) same-identity reconstruction and (b) cross-identity reenactment on the VoxCeleb1 and CelebV-HQ datasets. Our method can generate high-quality facial images even under large pose variations.
@@
FOMM [24] 53.97 22.96 0.0474 0.2200 1.4037 0.1509 78.15 20.92 0.0685 0.2925 3.6098 0.2955
LIA [31] 57.51 22.88 0.0488 0.2293 1.5395 0.1500 86.64 19.58 0.0830 0.3159 3.3632 0.3430
DaGAN [16] 51.55 22.92 0.0492 0.2251 1.5740 0.1652 99.84 20.49 0.0781 0.3209 7.6075 0.3301
MCNet [15] 51.45 24.59 0.0402 0.1996 1.2363 0.1254 78.33 22.20 0.0640 0.2732 4.1386 0.2903
MRFA [25] 48.49 25.26 0.0370 0.1872 1.1823 0.1188 75.73 22.41 0.0625 0.2670 3.7166 0.2527
AniPortrait [32] 52.65 20.15 0.0637 0.2767 2.6543 0.2623 61.56 19.71 0.0748 0.2878 2.2397 0.2739
FYE [19] 43.25 19.54 0.0714 0.2954 2.7071 0.2652 62.55 19.58 0.0802 0.3006 4.8637 0.3029
@@
Table 1: Quantitative comparison with state-of-the-art methods for same-identity reconstruction on VoxCeleb1 and CelebV-HQ datasets. Our results are the best on the Voxceleb1 dataset and competitive on the CelebV-HQ dataset.
We compare our method with a series of open-source state-of-the-art methods, including non-diffusion-based FOMM [24], LIA [31], DaGAN [16], MCNet [15] and MRFA [25], and diffusion-based AniPortrait [32] and Follow-Your-Emoji (FYE) [19].
Same-identity Reconstruction. To evaluate the performance on same-identity reconstruction, we use the first frame of each video as the source image and reconstruct the whole video. Tab. 1 presents the quantitative results for same-identity reconstruction. Our method outperforms the other methods almost on all metrics on VoxCeleb1 dataset and remains competitive when generalized to the more challenging CelebV-HQ. Compared with those unsupervised methods [24, 31, 16, 15, 25], our method achieves better motion estimation, e.g., our method gets the best AKD score on CelebV-HQ dataset. This result verifies the effectiveness of our designed multi-scale motion codebook and its generalizability. For the results of image quality (i.e., FID, PSNR, $\mathcal{L}_{1}$, LPIPS), our method outperforms other methods on VoxCeleb1 dataset, even those diffusion methods [32, 19] trained with larger-scale datasets. It indicates that our designed multi-scale appearance codebook is capable to compensate for the intermediate warped feature for better talking head video generation.We also show some qualitative results in Fig. 6a. Our method can generate plausible unseen facial regions (row 2) and handle large motion (row 4) with accuracy.
Cross-identity Reenactment. We conduct cross-identity reenactment experiments to validate our method. The qualitative results shown in Fig. 6b indicate the superiority of our method. Compared to non-diffusion-based MCNet [15] and MRFA [25], our method better preserves facial details (row 1 and 3), even under large motions (row 2 and 4). Diffusion-based FYE [19] often produces exaggerated expressions and struggles to imitate the driving expressions accurately, likely due to its reliance on landmark-based embeddings without explicitly modeling motion. These findings confirm the effectiveness of our framework. More experimental results are shown in the supplementary material.
@@ 4.3 Ablation Study
We perform ablation studies to assess the effectiveness of the learned multi-scale motion and appearance compensatory codebooks. The model variants in Tab. 2 are as follows: (i) “Baseline" is the model without any compensatory codebook. (ii) “Baseline+SMC" includes only a single-scale motion codebook, compensating for the initial motion flow only at scale $1$, and the compensated result is used to warp multi-scale features. (iii) “Baseline+MMC" includes a multi-scale motion codebook to compensate for the motion flows across scales. (iv) “Baseline+MMC+SAC" adds a single-scale appearance codebook to (iii), compensating only the warped feature at scale $1$. (v) “Baseline+MMC+MAC" is our full model with both multi-scale motion and appearance codebooks. We present the quantitative results in Tab. 2 and qualitative comparisons of (i), (iii), and (v) in Fig. 9.
Effect of Joint Codebook Learning. To evaluate the effectiveness of our jointly learned multi-scale motion and appearance codebooks, we visualize the reconstructed multi-scale motion flows with the motion codebook in Fig. 7 and appearance features with the appearance codebook in Fig. 8. For the motion flows, results in Fig. 7 are output of the motion flow decoder $\mathcal{D}_{M}$ given the quantized motion flow features. Despite the limited number of codes, our multi-scale motion codebook can reconstruct high-quality motion flows, confirming its ability to capture typical local motion patterns. For appearance features, we visualize the quantized features directly in Fig. 8. Despite some quantization loss, the multi-scale appearance codebook reconstructs the driving features well with limited codes, demonstrating its ability to store informative local appearance details. These results demonstrate the effectiveness of joint codebook learning, allowing the codebooks to store expressive local motion and appearance patterns.
@@
Figure 10: Visualization of the motion flow compensation process. We present the initial motion flow $\mathbf{M}^{0}$, the motion flow residuals $\{\mathbf{M}_{r}^{i}\}_{i=1}^{N}$ and the compensated motion flows $\{\mathbf{M}^{i}\}_{i=1}^{N}$.
Figure 11: Visualization of the appearance compensation results. We present the warped source features $\{\overline{\mathbf{F}}_{w}^{i}\}_{i=1}^{N}$ and the compensated appearance features $\{\overline{\mathbf{F}}_{c}^{i}\}_{i=1}^{N}$.
Effect of Multi-scale Appearance Codebook Compensation. The fourth row in Tab. 2 shows that adding single-scale appearance codebook compensation on top of multi-scale motion codebook compensation improves FID, LPIPS, and AKD, indicating that appearance codebook compensation refines warped source features for more realistic image generation. However, there is a slight drop in PSNR, $\mathcal{L}_{1}$, and AED, likely due to a conflict between scale $1$ compensated appearance features and other warped features with warping artifacts. The fifth row shows consistent improvement, suggesting that multi-scale appearance codebook compensation resolves this conflict and further refines warped features across scales, boosting overall performance. The last two columns in Fig. 9 also verify that multi-scale appearance codebook compensation leads to more accurate motion (e.g., the mouth in row 1, eyes in row 2, and shoulders in row 3) with realistic facial details (e.g., the hair in row 2). Additionally, Fig. 11 visualizes the compensated feature maps at different scales, showing more complete facial shapes and details compared to the warped feature $\overline{\mathbf{F}}_{w}^{i}$. These results validate the effectiveness of multi-scale appearance codebook compensation.
@@ 5 Conclusion
Acknowledgements. This research is supported in part by the Early Career Scheme of the Research Grants Council (RGC) of the Hong Kong SAR under grant No. 26202321, SAIL Research Project, HKUST-Zeekr Collaborative Research Fund, HKUST-WeBank Joint Lab Project, and Tencent Rhino-Bird Focused Research Program.
Supplementary Material
6 Additional Implementation Details
@@
We perform multi-scale compensation across $N=4$ scales. We employ the keypoint-based motion flow estimator from FOMM [24]. The multi-scale motion flows are estimated at a size of $64\times 64$. We use convolution layers to encode the motion flows into a latent motion flow space of size $32\times 32\times 32$ and set the multi-scale motion codebook size to $K=1024$ and $d_{m}=32$. We also use convolution layers to decode the quantized motion flow features while adopting the motion flow updater in MRFA [25] as our motion flow residual decoder. We employ the image encoder and decoder architecture from VQGAN [7] and further encode the multi-scale appearance features into a size of $32\times 32\times 256$. The multi-scale appearance codebook size is set to $T=1024$ and $d_{a}=256$.
@@
We follow the unsupervised training pipeline from FOMM [24], where the source and driving frames are extracted from the same video, and our framework learns to reconstruct the driving frame. For the training objective, we use the perceptual loss from FOMM [24] along with the L1 loss as the image reconstruction loss, and we set the loss weights as $\lambda_{adv}=0.8$, $\lambda^{1}=0.5$, $\lambda_{recon,m}=32$ and $\beta=0.25$. The entire framework is trained end-to-end utilizing the Adam optimizer, with a learning rate set to $8\times 10^{-5}$ and a batch size of 16 for 250K iterations on four NVIDIA RTX 3090 GPUs.
7 More Details on Experiments
7.1 Additional Details on the Compared Methods
@@
We evaluate the performance of the compared methods using their released pre-trained models, and we present the training datasets used for each method in Tab. 3. All the GAN-based methods [24, 31, 16, 15, 25] and our method are trained on the VoxCeleb1 [20] training set, while the diffusion-based methods AniPortrait [32] and Follow-Your-Emoji (FYE) [19] are trained on larger-scale datasets, including VFHQ [34], CelebV-HQ [46], HDTF [42], and their own collected dataset [19].
7.2 More Experimental Results
7.2.1 Video Results
We present video results for the ablation study and state-of-the-art comparisons on the project page [^1] https://shaelynz.github.io/synergize-motion-appearance/ to demonstrate the effectiveness of our video generation approach.
7.2.2 More Comparison Results
Cross-identity Reenactment. In the absence of ground truth for cross-identity reenactment, we conduct a user study comparing our approach to recent state-of-the-art methods, including a GAN-based model (MRFA [25]) and a diffusion-based model (Follow-You-Emoji (FYE) [19]). We randomly selected 10 source-driving pairs and asked 30 participants to evaluate the generated videos based on appearance realism, motion naturalness, and overall quality. The results shown in Fig. 12 indicate that users prefer our method, confirming its superiority.
Method Training Dataset FID $\downarrow$ CSIM $\uparrow$ ARD $\downarrow$
AniPortrait [32] VFHQ [34], CelebV-HQ [46] 66.61 0.7226 2.9146
FYE [19] HDTF [42], VFHQ [34], their collected dataset [19] 60.05 0.7558 3.0822
FOMM [24] VoxCeleb1 [20] 80.00 0.6010 1.8331
LIA [31] VoxCeleb1 [20] 72.55 0.6505 2.5404
DaGAN [16] VoxCeleb1 [20] 85.32 0.5743 2.0604
MCNet [15] VoxCeleb1 [20] 82.72 0.5618 1.6970
MRFA [25] VoxCeleb1 [20] 77.63 0.5962 1.5903
Ours VoxCeleb1 [20] 76.47 0.6142 1.6234
Table 3: Quantitative comparison for cross-identity reenactment on VoxCeleb1 dataset. AniPortrait [32] and Follow-Your-Emoji (FYE) [19] are trained on much larger-scale datasets and are not suitable for a direct comparison.